OpenStudy (stephaun):

Solve the system of equations. x + 3y = -1 2x + 2y = 6

OpenStudy (stephaun):

I really dont know how to do this

OpenStudy (anonymous):

does the question ask to solve both equation into one answer or solve each separately

OpenStudy (stephaun):

here is the set of answers

OpenStudy (anonymous):

ok

OpenStudy (stephaun):

(2, -1) (4, -1) (5, -2)

OpenStudy (stephaun):

and A was (-4, 1)

OpenStudy (texaschic101):

x + 3y = -1 ----> x = -3y - 1 now sub -3y - 1 in for x in the other equation 2x + 2y = 6 2(-3y - 1) + 2y = 6 -6y - 2 + 2y = 6 -6y + 2y = 6 + 2 -4y = 8 y = -2 now sub -2 in for y in either equation to find x x + 3y = -1 x + 3(-2) = -1 x - 6 = -1 x = -1 + 6 x = 5 solution is : (5,-2)

OpenStudy (stephaun):

?

OpenStudy (anonymous):

the answer is d

OpenStudy (stephaun):

Wow ima right this down for my notes thank you guys so much :3

OpenStudy (texaschic101):

anytime :)